# Berlin, Germany

### v2

@@ Abstract
We propose a natural intrinsic extension of the ridge regression from Euclidean spaces to general manifolds, which relies on Riemannian least-squares fitting, empirical covariance, and Mahalanobis distance. We utilize it for time-series prediction and apply the approach to forecast hurricane tracks and their wind speeds.
Keywords:
Hurricane Forecasting Mahalanobis Prediction Manifold-valued Tikhonov.
@@ 1 Introduction
Ridge regression, a well-established extension of Tikhonov regularization, is widely used in various fields such as computational science, engineering, and econometrics to estimate model parameters, especially in scenarios with highly correlated variables. It often improves prediction accuracy over simple linear regression by mitigating overfitting that arises from complex models [15, 12, 23]. However, many modern data types, such as directional data, shapes, or trajectories on the Earth’s surface, naturally reside on non-linear, curved spaces. Ignoring this inherent geometric structure by applying Euclidean techniques may lead to inaccurate models and suboptimal predictions. In this work, we propose a novel, intrinsic extension of ridge regression from traditional Euclidean spaces to arbitrary Riemannian manifolds. This extension leverages Riemannian least-squares fitting, empirical covariance, and the Mahalanobis distance framework, enabling robust analysis and prediction of manifold-valued time-series data. A core component of our approach is representing trajectories by best-fitting intrinsic Bézier curves, constructed via a natural generalization of the classic de Casteljau algorithm to manifold setting [20, 13]. This formulation extends geodesic regression to a polynomial regression paradigm suitable for curved spaces.
Our main contributions are threefold: 1. We formulate a natural, geometry-aware and intrinsic version of ridge regression that operates directly on manifolds, allowing for effective prediction of manifold-valued trajectories. 2. We derive explicit formulas for the gradient of the resulting objective functions, facilitating efficient numerical optimization. 3. We demonstrate the versatility of our method by applying it to the forecasting of hurricane tracks and their intensities, using real-world data.
@@
This work is organized as follows. The next section is devoted to mathematical preliminaries from Riemannian geometry, followed by a section on least-squares technique, the polynomial manifold-valued model, and our approach, the ridge regression. In Section 4, we discuss the application of our approach to hurricanes, certain modifications of the proposed approach, and numerical results.
2 Preliminaries
We provide a short intuitive background on Riemannian geometry and refer to [8] and [14] for this topic. A Riemannian manifold $M$ is a smooth space where each point has a smoothly varying inner product $g_{x}$ on its tangent space $T_{x}M$ for $x\in M$. Geodesics are curves that locally minimize distance, serving as the "straight lines" of curved spaces. The Riemannian exponential map at $x$ takes a tangent vector and maps it to a point in the manifold following the geodesic emanating from $x$. The local inverse, the logarithmic map, returns the initial velocity of a geodesic connecting $x$ to another point $y$ in $M$. Distances $dist(x,y)$ are calculated as the lengths of these geodesics. Jacobi fields are vector fields along a geodesic in a Riemannian manifold that describe how nearby geodesics diverge or converge. They arise from the second variation of the energy functional or length of geodesics and reflect the curvature of the manifold. Properties and computations that depend solely on the geometry of the manifold itself, not on an embedding space, are called intrinsic.
3 Prediction for Trajectories on Riemannian Manifolds
In this section, we provide an overview of our regression model to represent trajectories by best-fitting Bézier curves, and subsequently expose our approach to prediction.
@@ 3 Prediction for Trajectories on Riemannian Manifolds
Let $M$ be a finite-dimensional Riemannian manifold. We denote its Riemannian distance by $dist$, the exponential and its local inverse, the logarithmic maps of $M$ by $\exp$ and $\log$, respectively, and set $I:=[0,1]$. Moreover, we denote the tangent space of $M$ at $x$ by $T_{x}M$, and the number of control points of the best-fitting Bézier curves (polynomials) by $n$, and set $\mathcal{M}:=M^{n}$, $\mathcal{U}:=U^{n}$, where $U$ refers to a normal convex (also called totally normal) neighbourhood in $M$.
@@ 3.1 Bézier Polynomials and Regression
Let $b_{1},\cdots,b_{n}$ be points in $U$. Application the de Casteljau algorithm, defines a polynomial $p(t;b)$ with control points $b_{1},\cdots,b_{n}$. Here $b$ stands for $(b_{1},\cdots,b_{n})$ and $n-1$ is the degree of the polynomial $p(.;b)$. For a comprehensive introduction and some applications in data science, we refer to [16] and the recent overview [13]. Evaluation of the polynomials is, due to the use of the de Casteljau algorithm, computationally very efficient.
@@
We use the widespread least-squares regression technique to estimate the best-fitting (not restricted to interpolatory) Bézier curves that model the trajectories in $M$.
@@
Consider scalars $t_{1}<t_{2}<\cdots<t_{m}$ and sample points $y_{1},\cdots,y_{m}\in U$ of a trajectory $y$, i.e., $y_{i}=y(t_{i})$. Our regression model aims at finding a polynomial $p=p(.;b)$ that best fits the data $(t_{i},y_{i})$ in the least squares sense:
@@
The minimizer is the best-fitting polynomial of degree $n-1$ for data $(t_{i},y_{i})$ with $i=1,\cdots,m$. Applying a linear bijection, we may assume that $t_{1}=0$ and $t_{m}=1$. Computationally, we employ the parametrization by the control points to determine the corresponding polynomial $p(.;b^{\ast})$, where $b^{\ast}:=\operatorname{arg\,min}H$ and
$H(b):=\sum_{i=1}^{m}dist^{2}(y_{i},p(t_{i};b))$
The choice $b_{i}=\gamma(\frac{i-1}{n-1})$ with $i=1,\cdots,n$ and $\gamma(t):=\exp_{y_{1}}(t\log_{y_{1}}y_{m})$, serves as a convenient initial guess. The predictive power of the regression model can be measured by the coefficient of determination, denoted $R^{2}$. To compute it, let $H_{min}:=H(b^{\ast})$ and denote the minimum of
$G(x):=\sum_{i=1}^{m}dist^{2}(y_{i},x))$
by $G_{min}$. We recall that the minimizer of $G$ is the (Fréchet) mean of the points $y_{1},\cdots,y_{m}$. Now, the coefficient of determination reads
$R^{2}=1-\frac{H_{min}}{G_{min}}.$
Note that $\frac{1}{m}H_{min}$ and $\frac{1}{m}G_{min}$ are the unexplained and total variances, respectively.
@@ 3.2 Ridge Regression
Next, we propose our extended ridge regression. This approach, adapting a Tikhonov regularization to our parametric regression, allows prediction building on the polynomial model function. To this end, we utilize the notions of covariance and Mahalanobis distance, extended from Euclidean spaces to general manifolds (cf. the survey [21]). The regularizer has the geometric interpretation of Mahalanobis distance to the fitted prior data.
We denote the Riemannian metric of $\mathcal{M}$ by $g$, and its distance, exponential, and logarithmic maps by $\mathrm{d}$, $\mathrm{Exp}$ and $\mathrm{Log}$, respectively. Now, fix $\mu\in\mathcal{U}$, a positive definite selfadjoint endomorphism $S$ of $T_{\mu}\mathcal{M}$ and consider the function $F_{\mu,S}:\mathcal{U}\to\mathbb{R}$ defined by
$F_{\mu,S}(b)$ $=H(b)+g_{\mu}(S\mathrm{Log}_{\mu}b,\mathrm{Log}_{\mu}b).$
Clearly, we may write
$g_{\mu}(S\mathrm{Log}_{\mu}x,\mathrm{Log}_{\mu}x)=g_{\mu}(W\mathrm{Log}_{m}ux,W\mathrm{Log}_{\mu}x)$
where $W$ is a positive definite endomorphism satisfying $WW^{T}=S$ (in the Euclidean case represented by the Tikhonov matrix) on $T_{\mu}\mathcal{M}$. Now, fix $\lambda>0$. We seek to minimize $F_{\mu,\lambda S}$ with $\mu$ the expected value, $S$ the precision, i.e., inverse of the covariance $\Sigma$ of $b$, and a suitably chosen ridge parameter $\lambda$.
@@ 3.2.1 Linearization and Gradient
Let $\textnormal{grad}\,F_{\mu,S}$ denote the gradient of $F_{\mu,S}$. If $M$ is Euclidean, then $p(t;b)=Xb$ with a matrix $X$ determined by $(t_{1},\cdots,t_{m})$, and stacking $y_{1},\cdots,y_{m}$ as a vector $y$, a straightforward computation shows that
$\textnormal{grad}_{b}F_{\mu,S}=X^{T}(Xb-y)+S(b-\mu),$
implying the following well-known explicit expression for the minimizer of $F_{\mu,S}$
$b=(X^{T}X+S)^{-1}(X^{T}y+S\mu).$
Generally, due to absence of an explicit analytic solution, the regression task has to be solved numerically. To this end, we employ a Riemannian steepest descent solver [1]. In this regard, if we do not use automatic differentiation, the main challenge is to compute the gradient of $F_{\mu,S}$. It has been shown in [4, Sec. 4.2] that $\textnormal{grad}\,H$ can be computed as the adjoint of the sum of certain Jacobi fields (in closed form, if $M$ is a symmetric space). We still need to determine the gradient of the squared Mahalanobis distance. For this purpose, we have the following result.
Proposition 1
@@
Fix $x\in\mathcal{U}$. Let $\gamma$ be the geodesic given by $\gamma(t)=\mathrm{Exp}_{\mu}(t\mathrm{Log}_{\mu}x)$ with $t\in I$. Let $J$ be the Jacobi field along $\gamma$ with $J(0)=0$ and $\dot{J}(0)=S\mathrm{Log}_{\mu}x$ (dot stands for derivative with respect to $t$). Then
$\textnormal{grad}_{x}f=2J(1).$
Let $k$ be the function defined by $k(x)=\mathrm{Exp}_{\mu}(W\mathrm{Log}_{\mu}x)$ on $\mathcal{U}$, and denote the tangent map of $k$ evaluated at $x$ by $d_{x}k$. Then
$\textnormal{grad}_{x}f=-2(d_{x}k)^{*}(\mathrm{Log}_{k(x)}\mu),$
where the superscript * represents the adjoint.
@@
$=2g_{x}(d_{\mathrm{Log}_{\mu}x}\mathrm{Exp}_{\mu}(S\mathrm{Log}_{\mu}x),d_{\mathrm{Log}_{\mu}x}\mathrm{Exp}_{\mu}(d_{x}\mathrm{Log}_{\mu}v)),$
where the last equation follows from the Gauß lemma. Now, $\mathrm{Exp}_{\mu}(\mathrm{Log}_{\mu})$ is the identity on $\mathcal{U}$, hence we have $d_{\mathrm{Log}_{\mu}x}\mathrm{Exp}_{\mu}(d_{x}\mathrm{Log}_{\mu}v)=v$ and arrive at
$d_{x}fv$ $=2g_{x}(d_{\mathrm{Log}_{\mu}x}\mathrm{Exp}_{\mu}(S\mathrm{Log}_{\mu}x),v).$
@@
$\textnormal{grad}_{x}f=2d_{\mathrm{Log}_{\mu}x}\mathrm{Exp}_{\mu}(S\mathrm{Log}_{\mu}x).$
Now, let $\epsilon>0$ be small enough, such that the image of the map $C$ given by
$C(\tau,t):=\mathrm{Exp}_{\mu}(t(\mathrm{Log}_{\mu}x+\tau S\mathrm{Log}_{\mu}x)),\>(\tau,t)\in\,]-\epsilon,\epsilon[\times I,$
is in $\mathcal{U}$. The map $C$ is a variation of $\gamma$ through geodesics, and the vector field $J$ defined by $J(t)=\partial_{\tau}C(0,t)$ the Jacobi field along $\gamma$ with $J(0)=0$ and $\dot{J}(0)=S\mathrm{Log}_{\mu}x$. Due to [14, Sec. 5.2]
$d_{\mathrm{Log}_{\mu}x}\mathrm{Exp}_{\mu}(S\log_{\mu}x)=J(1),$
which immediately implies the first desired formula. Now, consider the function $r$ defined by $r=d^{2}(\mu,.)$ on $U$. We have $f(x)=r(k(x))$. Thus,
@@
$=g_{x}((d_{x}k)^{*}(\textnormal{grad}_{k(x)}r),v)$
and the well-known formula $\textnormal{grad}_{y}r=-2\mathrm{Log}_{y}\mu$ with $y=k(x)$ completes the proof.
We remark that if $M$ is embedded in some Euclidean space, then the gradient can also be obtained by projecting the Euclidean one to the tangent space at the point of the evaluation.

### v3

@@ Abstract
We propose a natural intrinsic extension of ridge regression from Euclidean spaces to general Riemannian manifolds for time-series prediction. Our approach combines Riemannian least-squares fitting via Bézier curves, empirical covariance on manifolds, and Mahalanobis distance regularization. A key technical contribution is an explicit formula for the gradient of the objective function using adjoint differentials, enabling efficient numerical optimization via Riemannian gradient descent. We validate our framework through synthetic spherical experiments (achieving significant error reduction over unregularized regression) and hurricane forecasting.
MSC 53C22 53A25 53B50 62M10
@@ 1 Introduction
Manifold-valued time series arise naturally across diverse scientific domains: directions, bird migration, hurricane tracks on the sphere, rigid body rotations in SO(3), diffusion tensor evolution on the manifold of symmetric positive-definite matrices, and shape trajectories, for instance in Kendall’s shape space. Traditional Euclidean methods for time-series analysis ignore the intrinsic geometric structure of these spaces, leading to artifacts such as non-unit quaternions when predicting rotations or non-positive-definite predictions in covariance modeling. Intrinsic methods that respect the geometry of data are essential for accurate and physically meaningful predictions.
We use intrinsically defined polynomials to model manifold-valued time-series. To this end, we utilize the extension of the ordinary de Casteljau algorithm and the resulting Bézier curves to manifolds, presented in [Popiel2007BzierCA] and [nava2013casteljau]. The latter also proposes some basic properties and presents applications to motion of rigid body, space of positive definite matrices, construction of canal and developable surfaces, and curve design on implicit surfaces and polyhedra. Moreover, the mentioned work also extends the ordinary de Boor algorithm, and therewith, Bézier splines (also called composite Bézier curves) to manifolds. Note that, variational splines [varsplines] also offer flexible models, but are computationally more costly. The work [singh2013hierarchical] characterizes manifold-valued polynomials by vanishing higher-order covariant derivative. However, this approach also suffers from high computational cost (polynomials are determined by solving higher-order differential equations involving the curvature tensor). Due to evaluation via de Casteljau algorithm, Bézier polynomials offer high computational efficiency.
We employ a natural extension of the widespread least-squares regression technique to estimate the best-fitting Bézier curves that model the trajectories. We recall that smoothing, dimensional reduction, and suppressing inconsistencies and noise are advantages of regression. For a comprehensive introduction and applications, we refer to [nava2019geomix, nava2020geo] and [nava2022HierCeoModel], which employed geodesic regression (as a counter part to the ordinary linear one) and [nava2023hur, haniknava2024dc] for the general case of manifold-valued splines. In the applications, along with the regression, we compute the coefficient of determination to examine the model’s adequacy. While geodesic regression extends linear regression to manifolds and variational splines [varsplines] offer flexible trajectory models, these methods lack regularization mechanisms for handling correlated or limited training data.
Ridge regression is a popular adaptation of Tikhonov regularization to standard regression to estimate coefficients in the underlying model, especially when independent variables are highly correlated. This method has been used in many fields, including computational science and engineering and econometrics. It often provides prediction efficiency over simple linear regression with a significant number of model parameters and helps reduce overfitting that can result from model complexity. For an overview, we refer to [overviewridgereg2009, golub1999tikhonov] and [overviewridge2022].
The main contributions of this work are as follows. We propose a natural intrinsic extension of the ordinary ridge regression from Euclidean spaces to arbitrary Riemannian manifolds, which incorporates the Mahalanobis distance and correlation matrix, and allows prediction of trajectories in the underlying manifold. We also derive an explicit formula for the gradient of the corresponding objective function, whose minimizer determines the model parameters. In addition, we validate our approach on synthetic spherical data, and hurricane tracks and intensities (wind speeds).
@@
This work is organized as follows. The next section is devoted to mathematical preliminaries from Riemannian geometry, least-squares technique, the polynomial manifold-valued model, and our approach, the ridge regression. In Section 3, we discuss applications including synthetic validation and hurricane forecasting, along with numerical results.
2 Prediction for Trajectories in Riemannian Manifolds
This section details our regression model for representing trajectories using best-fitting Bézier curves, followed by our approach to prediction. For background on Riemannian geometry, we refer the reader to [doCarmo1992] and the brief introduction in Appendix 5.1.
@@ 2 Prediction for Trajectories in Riemannian Manifolds
Let $M$ be a finite-dimensional Riemannian manifold with Riemannian distance distdist, exponential map $\exp$ and local inverse, logarithmic map $\log$, respectively, and set $I:=[0,1]$. Moreover, we denote the tangent space of $M$ in $x$ by $T_{x}M$, and the number of control points of the best-fitting Bézier curves (polynomials) by $n$, set $\mathcal{M}:=M^{n}$ (power manifold), and $\mathcal{U}:=U^{n}$, where $U$ refers to a normal convex neighborhood (also called totally normal) in $M$. Moreover, we set $\gamma(t;x,y):=\exp_{x}(t\log_{x}y)$ with $t\in I$ and $x,y\in U$.
@@ 2.1 Bézier Polynomials and Regression
To model manifold-valued time-series, we utilize the extension of the ordinary de Casteljau algorithm and the resulting Bézier curves to manifolds, presented in [Popiel2007BzierCA] and [nava2013casteljau]. The latter work not only introduces these geometric polynomials but also establishes fundamental properties and demonstrates applications across diverse domains: motion of rigid bodies in $\mathrm{SO}(3)$ , analysis in the space of positive definite matrices (crucial for diffusion tensor imaging and covariance modeling), construction of canal and developable surfaces in geometric design, and curve design on implicit surfaces and polyhedra. The versatility of this framework stems from its intrinsic formulation, which relies solely on the manifold’s geodesic structure rather than embedding-dependent constructions.
The work [singh2013hierarchical] characterizes manifold-valued polynomials by vanishing higher-order covariant derivatives, providing an alternative perspective rooted in differential geometry. However, this approach suffers from high computational cost, as polynomials are determined by solving higher-order differential equations involving the curvature tensor—a computationally intensive task, especially for manifolds with complex curvature structure. In contrast, due to evaluation via the de Casteljau algorithm, Bézier polynomials offer high computational efficiency through their recursive geometric construction, requiring only repeated geodesic interpolations without solving differential equations.
Moreover, we employ a natural extension of the widespread least-squares regression technique to estimate the best-fitting Bézier curves that model the trajectories. Regression techniques offer several crucial advantages: smoothing noisy observations, achieving dimensional reduction by representing complex trajectories through a small number of control points, and suppressing inconsistencies inherent in real-world measurements. These benefits are particularly valuable when dealing with heterogeneous data sources or measurements subject to varying levels of precision. For a comprehensive introduction and applications of regression on manifolds, we refer to [nava2019geomix, nava2020geo] and [nava2022HierCeoModel], which employed geodesic regression (as a counterpart to ordinary linear regression in Euclidean spaces) for shape analysis and epidemiological applications, and [nava2023hur, haniknava2024dc] for the general case of manifold-valued splines applied to trajectory analysis and morphometrics. In our applications, along with the regression, we calculate the determination coefficient ($R^{2}$) to examine the suitability of the model and ensure that the polynomial representation captures a substantial part of the variability in the observed trajectories.
@@
Let $b_{1},\cdots,b_{n}$ be points in $U$. These define by the de Casteljau algorithm a polynomial $p(t;b)$ with control points $b_{1},\cdots,b_{n}$. Here $b$ stands for $(b_{1},\cdots,b_{n})$ and n1n-1 is the degree of the polynomial $p(.;b)$. In the following, we briefly describe the extension of the de Casteljau algorithm and Bézier curves to Riemannian manifolds, and for a comprehensive introduction and some applications in data science, refer to [nava2023hur, haniknava2024dc].
Geometric Intuition. In Euclidean spaces, Bézier curves are constructed through repeated linear interpolation between control points. The elegance of the de Casteljau algorithm lies in its purely geometric construction: to evaluate a Bézier curve at parameter $t$, we recursively interpolate between consecutive control points, with each level of recursion reducing the number of points by one until a single point remains. This geometric construction generalizes naturally to Riemannian manifolds by replacing linear interpolation with geodesic interpolation.
2.1.1 De Casteljau Algorithm
For $t\in I$, we set
$p_{i}^{0}(t):=b_{i+1},$
$p_{i}^{r}(t):=\gamma(t;p_{i}^{r-1}(t),p_{i+1}^{r-1}(t)),\quad r=1,\dots,n-1,\quad i=0,\dots,n-r.$
We call $p(.;b):=p^{n-1}_{0}:I\to U$ Bézier curve of degree n1n-1 with control points $b_{1},\dots,b_{n}$. Figure 1 shows a visualization of the algorithm to construct a quadratic Bézier curve on the 2-sphere.
Figure 1: Construction of a quadratic Bézier curve on sphere via the de Casteljau algorithm. Control points $b_{1},b_{2},b_{3}$ (red) determine the curve (black). For parameter $t\in I$, we recursively interpolate: connect $b_{1},b_{2}$ and $b_{2},b_{3}$ with geodesics (blue), evaluate at $t$ to get intermediate points (green), then connect these with another geodesic and evaluate at t to obtain the curve point $p(t;b)$. This geometric construction generalizes naturally to higher degrees and arbitrary manifolds.
Computational Efficiency. Evaluation of Bézier polynomials is, due to the use of the de Casteljau algorithm, computationally very efficient. The algorithm requires $O(n^{2})$ geodesic interpolations to evaluate the curve at a single parameter value, where each interpolation involves computing the exponential and logarithmic maps. For symmetric spaces, including spheres, hyperbolic spaces, and spaces of symmetric positive definite matrices, these maps have closed-form expressions, making the evaluation extremely fast. Even for general manifolds where geodesics must be computed numerically, the recursive structure of the algorithm remains stable and efficient. Evaluation of Bézier polynomials is, due to the use of the de Casteljau algorithm, computationally very efficient.
2.1.2 Regression Model
@@
Next, we use the widespread least-squares regression technique to estimate the best-fitting (not restricted to interpolatory) Bézier curves that model the trajectories in $M$.
@@
Consider the scalars $t_{1}<t_{2}<\cdots<t_{m}$ and sample points $y_{1},\cdots,y_{m}\in U$ of a trajectory $y$, i.e., $y_{i}=y(t_{i})$. Our regression model aims to find a Bézier polynomial $p=p(.;b)$ that best fits the data $(t_{i},y_{i})$ in the least squares sense:
@@
The minimizer is the best-fitting polynomial of degree n1n-1 for data $(t_{i},y_{i})$ with $i=1,\cdots,m$. Applying a linear bijection, we may assume that $t_{1}=0$ and $t_{m}=1$. Computationally, we employ the parametrization by the control points to determine the corresponding polynomial $p(.;b^{\ast})$, where $b^{\ast}:=\operatorname{arg\,min}H$ and
$H(b):=\sum_{i=1}^{m}dist^{2}(y_{i},p(t_{i};b)),\>b\in\mathcal{U}$
The choice $b_{i}=\gamma(\frac{i-1}{n-1};y_{1},y_{m})$ with $i=1,\cdots,n$ serves as a convenient initial guess. We recall that the $R^{2}$ value (coefficient of determination)
$R^{2}=1-\frac{H_{\min}}{G_{\min}}$
measures the goodness of fit (closer to 1 = better fit), where
$G(x):=\sum_{i=1}^{m}dist^{2}(y_{i},x),\>x\in U.$
@@ 2.2 Ridge Regression
Next, we propose our extended ridge regression. This approach, adapting a Tikhonov regularization to our parametric regression, allows prediction building on the polynomial model function. To this end, we utilize the notions of covariance and Mahalanobis distance, extended from Euclidean spaces to general manifolds (cf. the survey [pennec2006intro]). The regularizer has the geometric interpretation of Mahalanobis distance to the fitted prior data.
We denote the Riemannian metric of $\mathcal{M}$ by $g$, and its distance, exponential, and logarithmic maps by $\mathrm{d}$, $\mathrm{Exp}$ and $\mathrm{Log}$, respectively. Now, fix $\lambda\geq 0$, $\mu\in\mathcal{U}$, let $S$ be a positive definite selfadjoint endomorphism of $T_{\mu}\mathcal{M}$ and consider the function $F_{\lambda}:\mathcal{U}\to\mathbb{R}$ defined by
$F_{\lambda}(b)$ $=H(b)+\lambda g_{\mu}(S\mathrm{Log}_{\mu}b,\mathrm{Log}_{\mu}b)$
$=\underbrace{\sum_{i=1}^{m}dist^{2}(y_{i},p(t_{i};b))}_{\text{data fidelity term}}+\lambda\underbrace{g_{\mu}(S\mathrm{Log}_{\mu}b,\mathrm{Log}_{\mu}b)}_{\text{regularization term}}.$
Clearly, we may write the regularization term as
$g_{\mu}(S\mathrm{Log}_{\mu}x,\mathrm{Log}_{\mu}x)=g_{\mu}(W\mathrm{Log}_{\mu}x,W\mathrm{Log}_{\mu}x)$
where $W$ is a positive definite endomorphism with $WW^{t}=S$ (e.g. the unique positive definite symmetric root of $S$) on $T_{\mu}\mathcal{M}$. $\sqrt{\lambda}W$ is in the Euclidean case called Tikhonov matrix. We seek to minimize $F_{\lambda}$ with $\mu$ the expected value, $S$ the precision, i.e., the inverse of the covariance $\Sigma$ of $b$, and a suitable ridge parameter $\lambda$.
@@ 2.2.1 Linearization and Gradient
Let $\textnormal{grad}\,F_{\lambda}$ denote the gradient of $F_{\lambda}$. If $M$ is Euclidean, then $p(t;b)=Xb$ with a matrix $X$ determined by $(t_{1},\cdots,t_{m})$, and stacking $y_{1},\cdots,y_{m}$ as a vector $y$, a straightforward computation shows that
$\textnormal{grad}_{b}F_{\lambda}=X^{T}(Xb-y)+\lambda S(b-\mu),$
implying the following well-known explicit expression for the minimizer of $F_{\lambda}$
$b=(X^{T}X+\lambda S)^{-1}(X^{T}y+\lambda S\mu).$
Generally, due to absence of an explicit analytic solution, the regression task has to be solved numerically. To this end, we employ a Riemannian steepest descent solver [riemopt]. In this regard, the main challenge is to calculate the gradient of $F_{\lambda}$. In [BergmanGousenbourger2018, Sec. 4.2] it has been shown that $\textnormal{grad}\,H$ can be computed as the adjoint of the sum of certain Jacobi fields (in closed form, if $M$ is a symmetric space). We still need to determine the gradient of the squared Mahalanobis distance. For this purpose, we have the following result.
Proposition 1(Gradient of squared Mahalanobis distance)
@@
Fix $x\in\mathcal{U}$. Let $\gamma$ be the geodesic given by $\gamma(t)=\mathrm{Exp}_{\mu}(t\mathrm{Log}_{\mu}x)$ with $t\in I$. Then
$\textnormal{grad}_{x}f=2(d_{x}\mathrm{Log}_{\mu})^{\dagger}S\mathrm{Log}_{\mu}x,$
where the superscript represents the adjoint. Moreover, let $k$ be the function defined by $k(x)=\mathrm{Exp}_{\mu}(W\mathrm{Log}_{\mu}x)$ on $\mathcal{U}$, and denote the tangent map of $k$ evaluated at $x$ by $d_{x}k$. Then
$\textnormal{grad}_{x}f=-2(d_{x}k)^{\dagger}\mathrm{Log}_{k(x)}\mu.$
@@
and arrive at
$d_{x}fv$ $=2g_{x}((d_{x}\mathrm{Log}_{\mu})^{\dagger}S\mathrm{Log}_{\mu}x,v).$
@@
$\textnormal{grad}_{x}f=2(d_{x}\mathrm{Log}_{\mu})^{\dagger}S\mathrm{Log}_{\mu}x.$
Now, consider the function $r$ defined by $r=d^{2}(\mu,.)$ on $U$. We have $f(x)=r(k(x))$. Thus,
@@
$=g_{x}((d_{x}k)^{\dagger}(\textnormal{grad}_{k(x)}r),v)$
The well-known formula $\textnormal{grad}_{y}r=-2\mathrm{Log}_{y}\mu$ with $y=k(x)$ completes the proof.
Note that we may compute the gradient using automatic differentiation tools such as PyTorch or JAX. Moreover, if $M$ is embedded in some Euclidean space, then the gradient can also be obtained simply by projecting the Euclidean one to the tangent space at the point of evaluation. Finally, recall that for symmetric spaces Riemannian building blocks, and particularly the logarithmic map, are available as closed form expressions, accelerating the computations. For the computation of the adjoint map, we refer to Appendix 5.2.